You spread 0.1 mL volume of a 10^(-5) dilution onto a nutrient agar plate. After 24 hours of incubation at 37°C, there were 223 colonies of bacteria on the plate.

A.) What is the original concentration (OCD) of bacteria in the stock sample this dilution came from? (2 points)

B.) Using the OCD value from part A, determine the number of colonies that would be expected to grow on a plate that is inoculated with 0.1 mL volume of 10^(-7) dilution from this same stock of bacteria.
Answer #1

A) In this case given data are:

  • The amount of bacterial culture plated = 0.1 ml
  • Dilution factor of the bacterial culture = 10-5
  • Number colonies yielded = 223 colonies.

Number of colonies plated /ml may be calculated as:

  • No. of colonies yielded when 0.1 ml of bacterial culture is plated =223

Therefore,

No. of colonies yielded when 1 ml of bacterial culture is plated = (223/0.1) x 1

= 2230

  • Colony forming unit per millilitre (CFU/ml) may be calculated as:

(CFU/ml) = Number of colonies per ml plated/ dilution factor

Therefore,

(CFU/ml) = 2200/10-5

= 2230 x 105

2.23 x 108 CFU/ml

B)

(CFU/ml) = Number of colonies per ml plated/ dilution x volume plated

For 10-7 dilution, and 0.1 ml plated:

2.23 x 108 CFU/ml = Number of colonies per ml plated / 10-7 x 0.1

Number of colonies per ml plated = 2.23 x 108 x 0.1/ 107= 2.23 =2 approximately